A unit cell consists of a cube in which there are anions(Y) at each corner and cations (X)at the centre of alteernate faces of the unit cell .What is the formula of compound?

A  unit cell consists of a cube in which there are anions(Y) at each corner and cations (X)at the centre of alteernate faces of the unit cell .What is the formula of compound?

Grade:12

1 Answers

Akil S
5 years ago
FCC - 
Contribution of an atom placed at corner of the unit cell is 1/8.
Therefore, the total no. of anion (B) which are present in a unit cell is = 1/8 × total no. of corners = 1/8 × 8 = 1 anion
Contribution of an atom placed at face centre = 1/2
When cations are placed at the alternative faces, the otal no. of face centred atoms are 2 as shown below in:
So, total no. of cations are = 1/2 × 2 = 1
Thus, the formula of compound is AB.